The probability that a given tourist goes to the amusement park is 0.47, and the probability that she goes to the water park is 0.58. If the probability that she goes either to the water park or the amusement park is 0.95, what is the probability that she visits both of the parks on vacation?
Answer by edjones(8007)

Let x=the probability that she goes to both parks.
.47+.58-x=.95
1.05-x=.95
-x=-.1
x=0.1
.
Ed